# Solar rotation and activity for cycle 24 from SDO/AIA observations

### v1

@@ 1 Introduction
Solar differential rotation may create magnetic activities resulting from the dynamo process (Babcock, 1961; Li et al., 2014). The differential rotations of the Sun cause the twisting of the magnetic field, which creates magnetic phenomena such as sunspots, flares, prominence, and brightenings in the solar atmosphere (Chowdhury et al., 2013; Sharma et al., 2020). Hence, measuring the variation trend of differential rotation can be important in understanding the dynamics inside the Sun and subphotospheric layer. Measuring the solar rotation profile parameters and the impact of rotation/differential rotation on solar activity is essential to understanding the dynamics of magnetic features (Clement, 1970; Vainshtein, 1976; Bratenahl et al., 1980; Ruediger, 1989). Investigating the rotation profile parameters within the activity cycles and their asymmetric characteristic within the activity cycles is critical in solar physics (Clark et al., 1979; Javaraiah et al., 2005; Vats & Chandra, 2011; Javaraiah, 2020).
@@
Here, we use CBPs as tracers to derive the solar differential rotation in 9 yr (2011--2019) of cycle 24 observed by SDO/AIA at 193 Å 4k (4096 $\times$ 4096) images. First, we applied a machine learning-based algorithm to identify and track CBP (Alipour & Safari, 2015; Hosseini Rad et al., 2020). The support vector machine (SVM) classifier uses Zernike moments (ZMs) of CBP and non-CBP features in this work. The ZMs contain unique information for each feature that can reconstruct the original image. The tracking algorithm is modified to distinguish two or more CBPs that emerged consecutively in a small box with a time interval of less than 10minutes. The present analysis considers the CBPs with at least 100minute duration to determine solar rotation parameters. Then, the central meridian distance (CMD) and latitude distance are selected. The sidereal and meridian velocities in the units per day are estimated for CBPs within the central solar equatorial regions by applying a least-squares fitting method. The monthly and annual values of rotational parameters for the area of interest in the solar surface and both hemispheres are determined. We investigate the asymmetry property of hemispheres from the solar activity and rotational parameters point of view and their correlations for cycle 24. Finally, we calculate the horizontal Reynolds stress as an indication of transferring momentum (angular) toward the equator.
@@ 3.1 Detection of CBPs
Here, we aim to identify and track CBPs to calculate the rotational velocity. CBPs appear in various structures, sizes, and durations. So, the traditional identification methods, e.g., threshold-based methods, faced several difficulties in identification and tracking. However, the recently developed machine learning algorithm was applied to identify and track brightening features from AIA and Solar Orbiter observations (Alipour et al., 2012; Javaherian et al., 2014; Alipour & Safari, 2015; Hosseini Rad et al., 2020; Shokri et al., 2022; Alipour et al., 2022).
@@ 4 Results and discussion
Figure 5 represents the variation of $\delta$ function (Equation 9) corresponding with different trail heights above the photosphere. The $\delta$ function indicates the differences in the central longitude sector rotation profile from neighboring sectors. The figure shows the parabola-like behavior for the $\delta$ function with a minimum at h = 5627 km. The minimum of $\delta$ implies that the slight differences in rotation profile (for different longitudes at the same latitude) originated from the invariance of solar rotation for different longitudes. The minimum value of the function at h = 5627 km implies that CBPs are situated around this height above the photosphere, as seen in 193 Å observations. As reported in the literature (Kwon et al., 2010; Sudar et al., 2016; Madjarska, 2019; Hosseini Rad et al., 2020), the CBPs mostly appeared 5000--10,000 km above the photosphere. Tian et al. (2007) emphasize that most CBPs observed from Fe xii passband are located below 20,000 km with an average height of 5000 km above the photosphere. Therefore, using Equations (6) and (7), we corrected latitude and the rotation velocity of CBPs for the height h = 5627 km, respectively.
@@
To investigate adequately and determine the dependencies of rotation parameters with solar activity (sunspots and flares), we used the scatter plot representation of the parameters. Figure 7 represents the scatter plot of $A$ and $B$ parameters versus yearly mean total sunspot numbers (top row) and flare index (bottom row). The slope of the fitted straight line (red line) is presented for each diagram. Considering the weights related to the errors of parameters $A$ and $B$, we obtained the slope of the fitted straight line for the scatter plot of rotation parameters and yearly mean total sunspot numbers/flare index for 9 yr of cycle 24. We observed that the slopes are significantly small, implying a fragile dependency on solar activity and rotation parameters within cycle 24 that does not agree with the theoretical model finding. The theoretical model predicted the negative correlation between the equatorial rotation parameter and solar activity (Brun, 2004; Lanza, 2006; Brun, 2014). A similar result was obtained by Poljančić Beljan et al. (2017, 2022). This discrepancy with the expected negative correlation of solar activity and parameter $A$ may be related to in-cycle variation due to the complex behavior of the Sun activity. The solar magnetic activity shows a secular increase for most years of the 20th century up to 1970 and then a decrease up to the end of cycle 24. Several studies support a long-term slowing of the Sun rotation through the last decades of the 20th century (Usoskin, 2017; Petrovay, 2020; Taran et al., 2022; Mohammadi Gouneh et al., 2023). Interestingly, Zhang et al. (2015) and Ruždjak et al. (2017) reported a temporary (secular) increase in rotation speed during the 1990s, followed by a return to the slowing trend.

### v2

@@ 1 Introduction
Solar differential rotation may create magnetic activities resulting from the dynamo process (Babcock, 1961; Li et al., 2014). The differential rotations of the Sun cause the twisting of the magnetic field, which creates magnetic phenomena such as sunspots, flares, prominence, and brightenings in the solar atmosphere (Chowdhury et al., 2013; Sharma et al., 2020). Hence, measuring the variation trend of differential rotation can be important in understanding the dynamics inside the Sun and subphotospheric layer. Measuring the solar rotation profile parameters and the impact of rotation/differential rotation on solar activity is essential to understanding the dynamics of magnetic features (Clement, 1970; Vainshtein, 1976; Bratenahl et al., 1980; Ruediger, 1989). Investigating the rotation profile parameters within the activity cycles and their asymmetric characteristic within the activity cycles is critical in solar physics (Clark et al., 1979; Javaraiah et al., 2005; Vats & Chandra, 2010; Javaraiah, 2020).
@@
Here, we use CBPs as tracers to derive the solar differential rotation in 9 yr (2011--2019) of cycle 24 observed by SDO/AIA at 193 Å 4k (4096 $\times$ 4096) images. First, we applied a machine learning-based algorithm to identify and track CBP (Alipour & Safari, 2015; Hosseini Rad et al., 2021). The support vector machine (SVM) classifier uses Zernike moments (ZMs) of CBP and non-CBP features in this work. The ZMs contain unique information for each feature that can reconstruct the original image. The tracking algorithm is modified to distinguish two or more CBPs that emerged consecutively in a small box with a time interval of less than 10minutes. The present analysis considers the CBPs with at least 100minute duration to determine solar rotation parameters. Then, the central meridian distance (CMD) and latitude distance are selected. The sidereal and meridian velocities in the units per day are estimated for CBPs within the central solar equatorial regions by applying a least-squares fitting method. The monthly and annual values of rotational parameters for the area of interest in the solar surface and both hemispheres are determined. We investigate the asymmetry property of hemispheres from the solar activity and rotational parameters point of view and their correlations for cycle 24. Finally, we calculate the horizontal Reynolds stress as an indication of transferring momentum (angular) toward the equator.
@@ 3.1 Detection of CBPs
Here, we aim to identify and track CBPs to calculate the rotational velocity. CBPs appear in various structures, sizes, and durations. So, the traditional identification methods, e.g., threshold-based methods, faced several difficulties in identification and tracking. However, the recently developed machine learning algorithm was applied to identify and track brightening features from AIA and Solar Orbiter observations (Alipour et al., 2012; Javaherian et al., 2014; Alipour & Safari, 2015; Hosseini Rad et al., 2021; Shokri et al., 2022; Alipour et al., 2022).
@@ 4 Results and discussion
Figure 5 represents the variation of $\delta$ function (Equation 9) corresponding with different trail heights above the photosphere. The $\delta$ function indicates the differences in the central longitude sector rotation profile from neighboring sectors. The figure shows the parabola-like behavior for the $\delta$ function with a minimum at h = 5627 km. The minimum of $\delta$ implies that the slight differences in rotation profile (for different longitudes at the same latitude) originated from the invariance of solar rotation for different longitudes. The minimum value of the function at h = 5627 km implies that CBPs are situated around this height above the photosphere, as seen in 193 Å observations. As reported in the literature (Kwon et al., 2010; Sudar et al., 2016; Madjarska, 2019; Hosseini Rad et al., 2021), the CBPs mostly appeared 5000--10,000 km above the photosphere. Tian et al. (2007) emphasize that most CBPs observed from Fe xii passband are located below 20,000 km with an average height of 5000 km above the photosphere. Therefore, using Equations (6) and (7), we corrected latitude and the rotation velocity of CBPs for the height h = 5627 km, respectively.
@@
To investigate adequately and determine the dependencies of rotation parameters with solar activity (sunspots and flares), we used the scatter plot representation of the parameters. Figure 7 represents the scatter plot of $A$ and $B$ parameters versus yearly mean total sunspot numbers (top row) and flare index (bottom row). The slope of the fitted straight line (red line) is presented for each diagram. Considering the weights related to the errors of parameters $A$ and $B$, we obtained the slope of the fitted straight line for the scatter plot of rotation parameters and yearly mean total sunspot numbers/flare index for 9 yr of cycle 24. We observed that the slopes are significantly small, implying a fragile dependency on solar activity and rotation parameters within cycle 24 that does not agree with the theoretical model finding. The theoretical model predicted the negative correlation between the equatorial rotation parameter and solar activity (Brun, 2004; Lanza, 2006; Brun, 2013). A similar result was obtained by Poljančić Beljan et al. (2017, 2022). This discrepancy with the expected negative correlation of solar activity and parameter $A$ may be related to in-cycle variation due to the complex behavior of the Sun activity. The solar magnetic activity shows a secular increase for most years of the 20th century up to 1970 and then a decrease up to the end of cycle 24. Several studies support a long-term slowing of the Sun rotation through the last decades of the 20th century (Usoskin, 2017; Petrovay, 2020; Taran et al., 2022; Mohammadi Gouneh et al., 2023). Interestingly, Zhang et al. (2015) and Ruždjak et al. (2017) reported a temporary (secular) increase in rotation speed during the 1990s, followed by a return to the slowing trend.